OpenStudy (anonymous):

For what value of x is the rational expression below equal to zero? (x-4)/(x-6)

jimthompson5910 (jim_thompson5910):

set it equal to zero, then solve for x (x-4)/(x-6) = 0 x-4 = 0*(x-6) x-4 = 0 x = ???

OpenStudy (anonymous):

so it would 4 then..

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

WooHoo I understand!!! thank you

jimthompson5910 (jim_thompson5910):

you're welcome